# ObjectAdd: Adding Objects into Image via a Training-Free Diffusion Modification Fashion

### v5

@@ III-C2 Attention Information Injection
Herein, we use a $max$ function to ensure that the result before softmax would not be less than 1 (mask value) for object generation in user-specified area.
@@ III-D1 Attention Refocusing
Our motive mainly originates from an in-depth analysis of the cross-attention map $\bar{\mathcal{C}}_{k,:}^{\gamma t}$ for the added object as visualized in Fig. 3a. In the middle diffusion steps typically $0.3\cdot T<t<0.5\cdot T$, we find the central area of $\mathcal{W}$ described object will be well focused in the layer of $(H_{\gamma},W_{\gamma})=(32,32)$. However, the edge area would be mixed up with the out-of-box area. Therefore, we consider refocusing the attention pixel, to adjust a user-drawn box, i.e., the mask $\mathcal{M}^{\gamma}$, into the object-centralized area ¯$\bar{\mathcal{M}}^{\gamma}$.
@@
Finally, in Fig. 3d, we refocus the user-drawn area by updating the binary mask as ¯$\bar{\mathcal{M}}^{\gamma}$ where ¯$\bar{\mathcal{M}}^{\gamma}_{i,j}=1$ if attention pixel ($i,j$) falls into object-centralized area and 0 otherwise. For better performance, in Fig. 3e, we apply mathematical morphology [26] to remove small objects and fill holes inside the masked area.
@@ III-D2 Object Expansion
Through attention refocusing, we tentatively shift attention from the user-drawn area $\mathcal{M}^{\gamma}\in\mathbb{R}^{H_{\gamma}\times W_{\gamma}}$ to the object-centralized area ¯$\bar{\mathcal{M}}^{\gamma}\in\mathbb{R}^{H_{\gamma}\times W_{\gamma}}$. Unluckily, ¯$\bar{\mathcal{M}}^{\gamma}$ fails to cover all pixels of the object since the cross-attention map helps seize object-centralized pixels but can’t well capture the edge area. For ease of better prompted image inpainting, we upsample ¯$\bar{\mathcal{M}}^{\gamma}$ as ¯$\bar{\mathcal{M}}\in\mathbb{R}^{H\times W}$ and further propose object expansion in the latent ¯$\bar{\mathcal{I}}^{t}\in\mathbb{R}^{H\times W\times C}$ space to better model the edge of added objects.
Given $\bar{\mathcal{M}}$ locates the approximate location of the object, we expand the boundary of $\bar{\mathcal{M}}$, as an agent to include the object edge. For every ¯$\bar{\mathcal{M}}_{i,j}$, we consider its 8-connected component set as $\mathcal{S}_{\bar{\mathcal{M}}_{i,j}}=\{\bar{\mathcal{M}}_{x,y}\}_{x\in\{i,i-1,i+1\},\,y\in\{j,j-1,j+1\}}$. Accordingly, we obtain the 8-neighborhood pixels of latent vector ¯$\bar{\mathcal{I}}^{t}_{i,j}\in\mathbb{R}^{C}$, expressed as $\mathcal{S}_{\bar{\mathcal{I}}_{i,j}}=\{\bar{\mathcal{I}}_{x,y}\}_{x\in\{i,i-1,i+1\},\,y\in\{j,j-1,j+1\}}$ where ¯$\bar{\mathcal{I}}^{t}_{x,y}\in\mathbb{R}^{C}$. Our object expansion firstly intends to find out the seed pixels that are conditioned on:
¯$\begin{split}\bar{\mathcal{M}}_{i,j}==1\quad\textrm{and}\quad 0\in\mathcal{S}_{\bar{\mathcal{M}}_{i,j}},\end{split}$ (12)
that is, ¯$\bar{\mathcal{I}}_{x,y}^{t}$ is considered as one seed pixel if it matches the object area and at least one of its neighbors falls out of the object area. Then, for each neighbor ¯$\bar{\mathcal{I}}_{x,y}$ outside object area with ¯$\bar{\mathcal{M}}_{x,y}==0$, it indicates a potential edge pixel. We compute its distance to the seed pixel and the average as:
@@
¯$\bar{\mathcal{M}}_{x,y}=\left\{\begin{array}[]{ll}1,\textrm{\; if $D(\bar{\mathcal{I}}_{x,y}|\mathcal{I}_{i,j}$},\mathcal{S}_{\mathcal{I}_{i,j}})<\mathcal{H}_{2},\\ \bar{\mathcal{M}}_{x,y},\textrm{\; otherwise}.\end{array}\right.$ (14)
where $\mathcal{H}_{2}$ is a user-specified parameter. The mask of out-of-object pixel ¯$\bar{\mathcal{I}}_{x,y}$ is flipped if ¯$\bar{\mathcal{I}}_{x,y}$ is close to the in-object pixel and the average of all neighbors. Every time after going through all seeds ¯$\bar{\mathcal{I}}_{i,j}$, we set those flipped ¯$\bar{\mathcal{I}}_{x,y}$ as new seeds. This process goes iterativly until no $D(\bar{\mathcal{I}}_{x,y}|\mathcal{I}_{i,j},\mathcal{S}_{\mathcal{I}_{i,j}})<\mathcal{H}_{2}$ found.
As results, we better locate the object pixels, and swap the rest part of the latent to get the optimized latent ¯$\bar{\mathcal{I}}^{t}$, which similar to Eq. (7) is formulated as:
¯$\bar{\mathcal{I}}^{t}=(1-\bar{\mathcal{M}})\odot{\mathcal{I}}^{t}+\bar{\mathcal{M}}\odot\bar{\mathcal{I}}^{t}.$ (15)
After this step, the ¯$\bar{\mathcal{I}}^{t}$ is now a combination of detected object and the non-object area of original prompt derived latent $\mathcal{I}^{t}$. We apply the above swapping at a random time step between $0.3\cdot T<t<0.5\cdot T$. Since we apply this operation in a time step far away from diffusion end, the robustness of diffusion model would make up the missed detected area like the smaller area in Fig. 3f and unnatural edges. Details will be ablated in Sec. IV-D.
@@ III-E Real Image Editing
~$\tilde{\mathcal{I}}^{0}=F(\tilde{\mathcal{Z}}_{T};\mathcal{W}).$ (16)
In this way, we ensure that the object in ~$\tilde{\mathcal{I}}^{0}$ is generated in the user-specified location after denoising, fulfilling the role of layout control methods discussed in Sec. III-C.
@@
¯$\bar{\mathcal{I}}^{\tilde{t}}=(1-\mathcal{M}^{\prime})\odot\bar{\mathcal{I}}^{\tilde{t}}+\mathcal{M}^{\prime}\odot\tilde{\mathcal{Z}}_{\tilde{t}}.$ (17)
@@ IV-A2 Compared Methods
Our ObjectAdd builds upon SD-v1-4 [5] as a pre-trained diffusion model, generating outputs at a $512\times 512$ resolution. As the first attempt to incorporate objects into diffusion-generated images, we compare our method with relevant works: DALL-E 3 [4], P2P [12], and SD-v1-4 [5]. Since these methods don’t cater to our task, we merge prompts $\mathcal{P}$ and $\mathcal{W}$ into a new prompt $\mathcal{P^{{}^{\prime}}}$. For example, if $\mathcal{P}=$ “A lake” and $\mathcal{W}=$ “A boat”, then $\mathcal{P^{{}^{\prime}}}=$ “A lake A boat”. Notably, P2P has a feature that highlights specific tokens, which we utilize to emphasize the second word-level prompt, i.e., $\mathcal{W}$.
@@ IV-A3 Evaluation Metrics
The first one is the “By Pixels”. We firstly remove the user-drawn area from both the original image, ${\mathcal{I}}^{0}$, and the edited image, ¯$\bar{\mathcal{I}}^{0}$. Then, we calculate the average absolute difference between each pixel of two images, formulated as:
$\mathcal{L}_{pixels}=avg(\lvert(1-\mathcal{M})\odot{\mathcal{I}}^{0}-(1-\mathcal{M})\odot\bar{\mathcal{I}}^{0}\rvert),$ (18)
@@
where the $clip()$ embeds the image and text into the same space, then calculates their distance. The edited images undergo masking to isolate the object. Our script for computing this metric is derived from the open-source codes [27].
We also apply FID [28] metric, since it is a commonly used one for measuring image generation quality.
@@ IV-F Limitation and Future Work
Moving forward, the focus will be on developing a refined approach that ensures seamless accessibility for non-specialists. Besides, exploring advanced open models like SD-v3 version [29] is crucial. Finally, creating a suite of automated evaluation tools for comprehensive assessment is of paramount importance. Besides, it would be an interesting idea to transform input image to prompt automatically using a pre-trained model like BLIP [30], which might make our method more practical for using.

### v6

@@ III-C2 Attention Information Injection
Herein, we use a maxmax function to ensure that the result before softmax would not be less than 1 (mask value) for object generation in user-specified area.
@@ III-D1 Attention Refocusing
Our motive mainly originates from an in-depth analysis of the cross-attention map $\bar{\mathcal{C}}_{k,:}^{\gamma t}$ for the added object as visualized in Fig. 3a. In the middle diffusion steps typically $0.3\cdot T<t<0.5\cdot T$, we find the central area of $\mathcal{W}$ described object will be well focused in the layer of $(H_{\gamma},W_{\gamma})=(32,32)$. However, the edge area would be mixed up with the out-of-box area. Therefore, we consider refocusing the attention pixel, to adjust a user-drawn box, i.e., the mask $\mathcal{M}^{\gamma}$, into the object-centralized area $\bar{\mathcal{M}}^{\gamma}$.
@@
Finally, in Fig. 3d, we refocus the user-drawn area by updating the binary mask as $\bar{\mathcal{M}}^{\gamma}$ where $\bar{\mathcal{M}}^{\gamma}_{i,j}=1$ if attention pixel ($i,j$) falls into object-centralized area and 0 otherwise. For better performance, in Fig. 3e, we apply mathematical morphology [soille1999morphological] to remove small objects and fill holes inside the masked area.
@@ III-D2 Object Expansion
Through attention refocusing, we tentatively shift attention from the user-drawn area $\mathcal{M}^{\gamma}\in\mathbb{R}^{H_{\gamma}\times W_{\gamma}}$ to the object-centralized area $\bar{\mathcal{M}}^{\gamma}\in\mathbb{R}^{H_{\gamma}\times W_{\gamma}}$. Unluckily, $\bar{\mathcal{M}}^{\gamma}$ fails to cover all pixels of the object since the cross-attention map helps seize object-centralized pixels but can’t well capture the edge area. For ease of better prompted image inpainting, we upsample $\bar{\mathcal{M}}^{\gamma}$ as $\bar{\mathcal{M}}\in\mathbb{R}^{H\times W}$ and further propose object expansion in the latent $\bar{\mathcal{I}}^{t}\in\mathbb{R}^{H\times W\times C}$ space to better model the edge of added objects.
Given $\bar{\mathcal{M}}$ locates the approximate location of the object, we expand the boundary of $\bar{\mathcal{M}}$, as an agent to include the object edge. For every $\bar{\mathcal{M}}_{i,j}$, we consider its 8-connected component set as $\mathcal{S}_{\bar{\mathcal{M}}_{i,j}}=\{\bar{\mathcal{M}}_{x,y}\}_{x\in\{i,i-1,i+1\},\,y\in\{j,j-1,j+1\}}$. Accordingly, we obtain the 8-neighborhood pixels of latent vector $\bar{\mathcal{I}}^{t}_{i,j}\in\mathbb{R}^{C}$, expressed as $\mathcal{S}_{\bar{\mathcal{I}}_{i,j}}=\{\bar{\mathcal{I}}_{x,y}\}_{x\in\{i,i-1,i+1\},\,y\in\{j,j-1,j+1\}}$ where $\bar{\mathcal{I}}^{t}_{x,y}\in\mathbb{R}^{C}$. Our object expansion firstly intends to find out the seed pixels that are conditioned on:
$\begin{split}\bar{\mathcal{M}}_{i,j}==1\quad\textrm{and}\quad 0\in\mathcal{S}_{\bar{\mathcal{M}}_{i,j}},\end{split}$ (12)
that is, $\bar{\mathcal{I}}_{x,y}^{t}$ is considered as one seed pixel if it matches the object area and at least one of its neighbors falls out of the object area. Then, for each neighbor $\bar{\mathcal{I}}_{x,y}$ outside object area with $\bar{\mathcal{M}}_{x,y}==0$, it indicates a potential edge pixel. We compute its distance to the seed pixel and the average as:
@@
$\bar{\mathcal{M}}_{x,y}=\left\{\begin{array}[]{ll}1,\textrm{\; if $D(\bar{\mathcal{I}}_{x,y}|\mathcal{I}_{i,j}$},\mathcal{S}_{\mathcal{I}_{i,j}})<\mathcal{H}_{2},\\ \bar{\mathcal{M}}_{x,y},\textrm{\; otherwise}.\end{array}\right$. (14)
where $\mathcal{H}_{2}$ is a user-specified parameter. The mask of out-of-object pixel $\bar{\mathcal{I}}_{x,y}$ is flipped if $\bar{\mathcal{I}}_{x,y}$ is close to the in-object pixel and the average of all neighbors. Every time after going through all seeds $\bar{\mathcal{I}}_{i,j}$, we set those flipped $\bar{\mathcal{I}}_{x,y}$ as new seeds. This process goes iterativly until no $D(\bar{\mathcal{I}}_{x,y}|\mathcal{I}_{i,j},\mathcal{S}_{\mathcal{I}_{i,j}})<\mathcal{H}_{2}$ found.
As results, we better locate the object pixels, and swap the rest part of the latent to get the optimized latent $\bar{\mathcal{I}}^{t}$, which similar to Eq. (7) is formulated as:
$\bar{\mathcal{I}}^{t}=(1-\bar{\mathcal{M}})\odot{\mathcal{I}}^{t}+\bar{\mathcal{M}}\odot\bar{\mathcal{I}}^{t}.$ (15)
After this step, the $\bar{\mathcal{I}}^{t}$ is now a combination of detected object and the non-object area of original prompt derived latent $\mathcal{I}^{t}$. We apply the above swapping at a random time step between $0.3\cdot T<t<0.5\cdot T$. Since we apply this operation in a time step far away from diffusion end, the robustness of diffusion model would make up the missed detected area like the smaller area in Fig. 3f and unnatural edges. Details will be ablated in Sec. IV-D.
@@ III-E Real Image Editing
$\tilde{\mathcal{I}}^{0}=F(\tilde{\mathcal{Z}}_{T};\mathcal{W}).$ (16)
In this way, we ensure that the object in $\tilde{\mathcal{I}}^{0}$ is generated in the user-specified location after denoising, fulfilling the role of layout control methods discussed in Sec. III-C.
@@
$\bar{\mathcal{I}}^{\tilde{t}}=(1-\mathcal{M}^{\prime})\odot\bar{\mathcal{I}}^{\tilde{t}}+\mathcal{M}^{\prime}\odot\tilde{\mathcal{Z}}_{\tilde{t}}$. (17)
@@ IV-A2 Compared Methods
Our ObjectAdd builds upon SD-v1-4 [rombach2022high] as a pre-trained diffusion model, generating outputs at a $512\times 512$ resolution. As the first attempt to incorporate objects into diffusion-generated images, we compare our method with relevant works: DALL-E 3 [betker2023improving], P2P [hertz2022prompt], and SD-v1-4 [rombach2022high]. Since these methods don’t cater to our task, we merge prompts $\mathcal{P}$ and $\mathcal{W}$ into a new prompt $\mathcal{P^{{}^{\prime}}}$. For example, if $\mathcal{P}=$ “A lake” and $\mathcal{W}=$ “A boat”, then $\mathcal{P^{{}^{\prime}}}=$ “A lake A boat”. Notably, P2P has a feature that highlights specific tokens, which we utilize to emphasize the second word-level prompt, i.e., $\mathcal{W}$.
@@ IV-A3 Evaluation Metrics
The first one is the “By Pixels”. We firstly remove the user-drawn area from both the original image, ${\mathcal{I}}^{0}$, and the edited image, $\bar{\mathcal{I}}^{0}$. Then, we calculate the average absolute difference between each pixel of two images, formulated as:
[…]
@@
where the $clip()$ embeds the image and text into the same space, then calculates their distance. The edited images undergo masking to isolate the object. Our script for computing this metric is derived from the open-source codes [taited2023CLIPScore].
We also apply FID [heusel2017gans] metric, since it is a commonly used one for measuring image generation quality.
@@ IV-F Limitation and Future Work
Moving forward, the focus will be on developing a refined approach that ensures seamless accessibility for non-specialists. Besides, exploring advanced open models like SD-v3 version [esser2024scaling] is crucial. Finally, creating a suite of automated evaluation tools for comprehensive assessment is of paramount importance. Besides, it would be an interesting idea to transform input image to prompt automatically using a pre-trained model like BLIP [li2022blip], which might make our method more practical for using.